# Leveraging Skills from Unlabeled Prior Data for Efficient Online Exploration

### v1

@@ 1 Introduction
The entanglement problem can be alleviated through hierarchical decomposition. Specifically, trajectories can broken into segments of task-agnostic skills, which are composed in various ways to solve various objectives. We posit that unlabeled trajectories thus present a twofold benefit, (1) as a way to learn a diverse set of skills, and (2) as off-policy examples of composing such skills. Notably, prior online RL methods that leverage pretrained skills largely ignore the second benefit, and discard the prior trajectories after the skills are learned (Ajay et al., 2021; Pertsch et al., 2021; Hu et al., 2023; Chen et al., 2024). We instead argue that such trajectories are critical, and can greatly speed up learning. We make use of a simple strategy of learning an optimistic reward model from online samples, and pseudo-relabeling past trajectories with an optimistic reward estimate. The past trajectories can thus be readily utilized as off-policy data, allowing for quick learning even with a very small number of online interactions.
@@ 2 Related Work
Unsupervised skill discovery. Unsupervised skill discovery methods first began in the online setting, where RL agents were tasked with learning structured behaviors in the absence of reward signal  (Gregor et al., 2016; Bacon et al., 2017; Florensa et al., 2017; Achiam et al., 2018; Eysenbach et al., 2018; Sharma et al., 2020; Hansen et al., 2020; Liu & Abbeel, 2021; Park et al., 2023b). These insights naturally transferred to the offline setting as a method of dealing with unlabeled trajectory data. Offline skill discovery methods largely comprise of two categories, those who extract skills based on optimizing unsupervised reward signals (in either the form of policies (Touati et al., 2022; Hu et al., 2023; Frans et al., 2024; Park et al., 2024) or Q-functions (Chen et al., 2024)), and those who utilize conditional behavior-cloning over subsets of trajectories  (Shankar & Gupta, 2020; Ajay et al., 2021; Singh et al., 2021; Pertsch et al., 2021; Nasiriany et al., 2022). Closest to our method in implementation are Ajay et al. (2021) and Pertsch et al. (2021), who utilize a trajectory-segment VAE to learn low-level skills, and learn a high-level policy online. However, in contrast to prior methods which all utilize offline data purely for skill-learning and do not keep it around during online training, we show that utilizing the data via relabeling is critical for fast exploration.
@@
Offline to online reinforcement learning. The offline-to-online reinforcement learning methods (Xie et al., 2021; Song et al., 2023; Lee et al., 2022; Agarwal et al., 2022; Zhang et al., 2023; Zheng et al., 2023; Ball et al., 2023; Nakamoto et al., 2024) focus on efficient online learning with the presence of offline data (often labeled with the reward value). Many offline RL approaches can be applied to this setting – simply run offline RL first on the offline data to convergence as an initialization and then continue training for online learning (using the combined dataset that consists of both offline and online data) (Kumar et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2024). However, such approaches often result in slow online improvements as offline RL objectives tend to overly constrain the policy behaviors to be close to the prior data, limiting the exploration capability. On the other hand, off-policy online RL methods can also be directly applied in this setting by directly treating the offline data as additional off-policy data in the replay buffer and learning the policy from scratch (Lee et al., 2022; Song et al., 2023; Ball et al., 2023). While related in spirit, these methods cannot be directly used in our setting as they require offline data to have reward labels.
@@
Data-driven exploration. A common approach for online exploration is to augment reward bonuses to the perceived rewards and optimize the RL agent with respect to the augmented rewards (Stadie et al., 2015; Bellemare et al., 2016; Houthooft et al., 2016; Pathak et al., 2017; Tang et al., 2017; Ostrovski et al., 2017; Achiam & Sastry, 2017; Burda et al., 2018; Ermolov & Sebe, 2020; Guo et al., 2022; Lobel et al., 2023). While most exploration methods operate in the purely online setting and focus on adding bonuses to the online replay buffer, recent works also start to explore a more data-driven approach that makes use of an unlabeled prior data to guide online exploration. Li et al. (2024) explore adding bonuses to the offline data, allowing them to optimize the RL agent to be optimistic about states in the data, encouraging exploration around the offline data distribution. Our method explores a similar idea of adding bonuses to the offline data but for training a high-level policy, allowing us to compose pretrained skills effectively for exploration. Hu et al. (2023) explore a slightly different strategy of learning a number of policies using offline RL that each optimizes for a random reward function. Then, it samples actions from these policies online to form an action pool from which the online agent can choose to select for exploration. This approach does not utilize offline data during the online phase and require all the policies (for every random reward function) to be represented separately. In contrast, our method makes use of the offline data as off-policy data for updating the high-level policy and our skills are represented using a single network (with the skill latent being the input to our network). As we will show in our experiments, being able to use offline data is crucial for learning to explore in the environment efficiently.
@@
Options framework. Our method is also related to the option framework (Sutton et al., 1999; Menache et al., 2002; Mannor et al., 2004; Şimşek & Barto, 2004; 2007; Konidaris, 2011; Daniel et al., 2016a; Srinivas et al., 2016; Daniel et al., 2016b; Fox et al., 2017; Bacon et al., 2017; Kim et al., 2019; Bagaria & Konidaris, 2019; Bagaria et al., 2024). Different from the approach we take that learns latent skills with a fixed time horizon ($H=4$ in all our experiments), the options framework provides a more flexible way to learn skills with varying time horizon, often defined by learnable initiation and/or termination conditions (Sutton et al., 1999). We opt for the simplified skill definition because it allows us to bypass the need to learn initiation or termination conditions, and frame the skill pretraining phase as a simple supervised learning task.
@@ 4 Skills from Unlabeled Prior data for Exploration (SUPE)
Practical implementation details. Following prior work on trajectory-segment VAEs (Ajay et al., 2021; Pertsch et al., 2021), we use a Gaussian distribution (with both mean and diagonal covariance learnable) for the trajectory encoder, the skill policy, as well as the state-dependent prior. While Pertsch et al. (2021) use a KL constraint between the high-level policy and the state-dependent prior, we use a simpler design without the KL constraint that works much better (as we show in Appendix E). To achieve this, we adapt the policy parameterization from (Haarnoja et al., 2018), where the action value is enforced to be between $-1$ and $1$ using a $\mathrm{tanh}$ transformation, and entropy regularization is applied on the squashed space. We use this policy parameterization for the high-level policy $\pi(z|s)$ to predict the skill action in the squashed space $z_{\mathrm{sqaushed}}$. We then recover the actual skill action vector by unsquashing it according to $z=\mathrm{arctanh}(z_{\mathrm{sqaushed}})$, so that it can be used by our skill policy $\pi_{\theta}(a|s,z)$. For upper-confidence bound (optimistic) estimation of the reward, ($r_{\mathrm{UCB}}(s_{0},\hat{z})$), we directly borrow the UCB estimation implementation in Li et al. (2024) (Section 3, practical implementation section in their paper), where they use a combination of the random network distillation (RND) (Burda et al., 2018) reward bonus and the predicted reward from a reward model (see Appendix C, Ours for more details). For the off-policy high-level agent, we follow Li et al. (2024) to use RLPD (Ball et al., 2023) that takes a balanced number of samples from the prior data and the online replay buffer for agent optimization. In addition to using the optimistic offline reward label, we also find that adding the RND reward bonus to the online batch is also helpful to encourage online exploration, so we use it in all our experiments.
@@ 5 Experimental Results
    Does offline data help our method to compose skills better for faster exploration?
(b) b) Kitchen
(c) c) Visual AntMaze: maze layout and image observation examples
Figure 2: We experiment on three challenging, sparse-reward domains: AntMaze, Kitchen, and Visual AntMaze. a): AntMaze (Fu et al., 2020) (state-based) with three different maze layouts (antmaze-medium, antmaze-large, and antmaze-ultra) and the corresponding four goal locations (denoted as the red dots) that we experiment with for each of the layouts; b): Kitchen (Fu et al., 2020) (state-based); c): Visual AntMaze (Park et al., 2023a) with colors added to the floor with local $64\times 64$ image observations (e.g., see examples right of the maze). The color of the floor uniquely identifies the ant’s position within the maze. For both state-based and visual AntMaze, the ant starts at the bottom-left corner in the beginning of every episode.
@@ 5.1 Experimental setup
We conduct our experiments on three challenging sparse reward domains (Figure 2).
Aside from the two state-based domains above, we also consider a visual domain below to test the ability of our method in scaling up to high-dimensional image observations.
We use the normalized return, a standard metric for D4RL (Fu et al., 2020) environments, as the main evaluation metric. For the Kitchen domain, the normalized return represents the average percentage of the tasks that are solved. The normalized return for the AntMaze domains (state-based and visual) represents the average success rate of reaching the goal.
@@ 5.2 Comparisons
While there is no existing method in our problem setting that utilizes unlabeled prior data in both the pretraining phase and the online learning phase, there are methods that make use of the prior data in either phase. We first consider two baselines that do not use pretraining and instead directly perform online learning.
Online. This baseline completely discards the offline data and the exploration is done with online reward bonus implemented by random network distillation (RND) (Burda et al., 2018). For all the baselines below as well as our method, we add online RND bonus to the replay buffer to encourage exploration.
ExPLORe (Li et al., 2024). This baseline is similar to our method in the sense that it also uses exploration bonus and offline data to encourage exploration. The one crucial difference is that it does not perform unsupervised skill pretraining and learns a 1-step policy directly online. As we will show, pretraining is crucial for our method to find goals faster and lead to more efficient online learning. It is worth noting that the original ExPLORe method does not make use of online RND. To make the comparison fair, we additionally add online RND bonus to this baseline to help it explore better online. For completeness, we also include the performance of the original ExPLORe method in the Appendix (Figure 8).
@@
Online with Skills. We also consider two skill-based baselines where the prior data is discarded in the online phase and the high-level policy is trained from scratch online with exploration bonus. We experiment with two types of pretraining skills. The first one is the trajectory VAE skill used in our method. The second one is from a recently proposed unsupervised offline skill discovery method where skills are pretrained to be able to traverse a learned Hilbert representation space (Park et al., 2024) (HILP). We use the exact same high-level RL agent as our method except that the agent no longer makes use of the prior data online.
Finally, we also consider a novel baseline that also uses prior data during pretraining and online exploration, but uses HILP skills rather than trajectory-based skills.
@@
For the visual antmaze environment, we use the same image encoder used in RLPD (Ball et al., 2023). We also follow one of our baselines, ExPLORe (Li et al., 2024), to use ICVF (Ghosh et al., 2023), a method that uses task-agnostic value functions to learn image/state representations from passive data. ICVF takes in an offline unlabeled trajectory dataset with image observations and pretrain an image encoder in an unsupervised manner. Following ExPLORe, we take the weights of the image encoder from ICVF pretraining to initialize the image encoder’s weights in the RND network. To make the comparison fair, we also apply ICVF to all our baselines (see details in Appendix D).
Figure 3: Aggregated normalized return across three different domains. Ours achieves the best performance through training on all three domains. ExPLORe achieves strong later stage performance on AntMaze, but struggles in high-dimensional Visual AntMaze and Kitchen tasks. Online w/ HILP Skills and HILP w/ Offline Data achieve decent initial return on Kitchen, but struggle to learn in all three domains. Online w/ Trajectory Skills consistently underperforms Ours across all three environments. Diffusion BC + JSRL learns reasonably well in Kitchen, but performs much worse in AntMaze and Visual AntMaze. Online does not perform competitively at any stage of exploration. Section 5.2 contains details on the baselines we compare with. Each curve is an average over 8 seeds. For AntMaze, we aggregate over 3 maze layouts and 4 goals. For Kitchen, we aggregate over 3 tasks. For Visual AntMaze, we aggregate over 4 goals on one maze layout.

### v2

@@ 1 Introduction
The entanglement problem can be alleviated through hierarchical decomposition. Specifically, trajectories can be broken into segments of task-agnostic skills, which are composed in various ways to solve various objectives. We posit that unlabeled trajectories thus present a twofold benefit, (1) as a way to learn a diverse set of skills, and (2) as off-policy examples of composing such skills. Notably, prior online RL methods that leverage pretrained skills largely ignore the second benefit, and discard the prior trajectories after the skills are learned (Ajay et al., 2021; Pertsch et al., 2021; Hu et al., 2023; Chen et al., 2024). We instead argue that such trajectories are critical, and can greatly speed up learning. We make use of a simple strategy of learning an optimistic reward model from online samples, and pseudo-relabeling past trajectories with an optimistic reward estimate. The past trajectories can thus be readily utilized as off-policy data, allowing for quick learning even with a very small number of online interactions.
@@ 2 Related Work
Unsupervised skill discovery. Unsupervised skill discovery methods first began in the online setting, where RL agents were tasked with learning structured behaviors in the absence of reward signal (Gregor et al., 2016; Bacon et al., 2017; Florensa et al., 2017; Achiam et al., 2018; Eysenbach et al., 2018; Sharma et al., 2020; Hansen et al., 2020; Park et al., 2023b). These insights naturally transferred to the offline setting as a method of dealing with unlabeled trajectory data. Offline skill discovery methods largely comprise of two categories, those who extract skills based on optimizing unsupervised reward signals (in either the form of policies (Touati et al., 2022; Hu et al., 2023; Frans et al., 2024; Park et al., 2024b) or Q-functions (Chen et al., 2024)), and those who utilize conditional behavior-cloning over subsets of trajectories (Paraschos et al., 2013; Merel et al., 2018; Shankar & Gupta, 2020; Ajay et al., 2021; Singh et al., 2021; Pertsch et al., 2021; Nasiriany et al., 2022). Closest to our method in implementation are Ajay et al. (2021) and Pertsch et al. (2021), who utilize a trajectory-segment VAE to learn low-level skills, and learn a high-level policy online. In contrast to prior methods which all utilize offline data purely for skill-learning and do not keep it around during online training, we show that utilizing the data via relabeling is critical for fast exploration.
@@
Offline to online reinforcement learning. The offline-to-online reinforcement learning methods (Xie et al., 2021b; Song et al., 2023; Lee et al., 2022; Agarwal et al., 2022; Zhang et al., 2023; Zheng et al., 2023; Ball et al., 2023; Nakamoto et al., 2024) focus on efficient online learning with the presence of offline data (often labeled with the reward value). Many offline RL approaches can be applied to this setting – simply run offline RL first on the offline data to convergence as an initialization and then continue training for online learning (using the combined dataset that consists of both offline and online data) (Kumar et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2024). However, such approaches often result in slow online improvements as offline RL objectives tend to overly constrain the policy behaviors to be close to the prior data, limiting the exploration capability. On the other hand, off-policy online RL methods can also be directly applied in this setting by directly treating the offline data as additional off-policy data in the replay buffer and learning the policy from scratch (Lee et al., 2022; Song et al., 2023; Ball et al., 2023). While related in spirit, these methods cannot be directly used in our setting as they require offline data to have reward labels.
@@
Data-driven exploration. A common approach for online exploration is to augment reward bonuses to the perceived rewards and optimize the RL agent with respect to the augmented rewards (Stadie et al., 2015; Bellemare et al., 2016; Houthooft et al., 2016; Pathak et al., 2017; Tang et al., 2017; Ostrovski et al., 2017; Achiam & Sastry, 2017; Merel et al., 2018; Burda et al., 2018; Ermolov & Sebe, 2020; Guo et al., 2022; Lobel et al., 2023). While most exploration methods operate in the purely online setting and focus on adding bonuses to the online replay buffer, recent works also start to explore a more data-driven approach that makes use of an unlabeled prior data to guide online exploration. Li et al. (2024) explore adding bonuses to the offline data, allowing them to optimize the RL agent to be optimistic about states in the data, encouraging exploration around the offline data distribution. Our method explores a similar idea of adding bonuses to the offline data but for training a high-level policy, allowing us to compose pretrained skills effectively for exploration. Hu et al. (2023) explore a slightly different strategy of learning a number of policies using offline RL that each optimizes for a random reward function. Then, it samples actions from these policies online to form an action pool from which the online agent can choose to select for exploration. This approach does not utilize offline data during the online phase and require all the policies (for every random reward function) to be represented separately. In contrast, our method makes use of the offline data as off-policy data for updating the high-level policy and our skills are represented using a single network (with the skill latent being the input to our network). As we will show in our experiments, being able to use offline data is crucial for learning to explore in the environment efficiently.
Hierarchical reinforcement learning. The ability of RL agents to explore and behave effectively over a long horizon is an important research goal in the field of hierarchical RL (HRL) (Dayan & Hinton, 1992; Dietterich, 2000; Vezhnevets et al., 2016; Daniel et al., 2016a; Kulkarni et al., 2016; Vezhnevets et al., 2017; Peng et al., 2017; Riedmiller et al., 2018; Nachum et al., 2018; Ajay et al., 2021; Shankar & Gupta, 2020; Pertsch et al., 2021; Gehring et al., 2021; Xie et al., 2021a). HRL methods typically learn a high-level policy to leverage a space of low-level primitive policies online. These primitives can be either manually specified (Dalal et al., 2021) or pre-trained using unsupervised skill discovery methods as discussed above. While many existing works learn or fine-tune the primitives online along with the high-level policy (Dietterich, 2000; Kulkarni et al., 2016; Vezhnevets et al., 2016; 2017; Nachum et al., 2018; Shankar & Gupta, 2020), others opt for a less flexible but simpler formulation where the primitives are kept fixed after an initial pre-training phase and only the high-level policy is being learned online (Peng et al., 2017; Riedmiller et al., 2018; Ajay et al., 2021; Pertsch et al., 2021; Gehring et al., 2021). Our work adopts the later strategy where we offline pre-train skills using a static, unlabeled dataset. None of prior HRL methods simultaneously leverage offline data for skill pre-training and as additional off-policy data for high-level policy learning online. As we show in our experiments, both of them are crucial in achieving sample efficient learning on challenging sparse-reward tasks.
@@
Options framework. Many existing works on building hierarchical agents also adopt the options framework (Sutton et al., 1999; Menache et al., 2002; Chentanez et al., 2004; Mannor et al., 2004; Şimşek & Barto, 2004; 2007; Konidaris, 2011; Daniel et al., 2016a; Srinivas et al., 2016; Daniel et al., 2016b; Fox et al., 2017; Bacon et al., 2017; Kim et al., 2019; Bagaria & Konidaris, 2019; Bagaria et al., 2024). Different from the approach we take that learns latent skills with a fixed time horizon ($H=4$ in all our experiments), the options framework provides a more flexible way to learn skills with varying time horizon, often defined by learnable initiation and/or termination conditions (Sutton et al., 1999). We opt for the simplified skill definition because it allows us to bypass the need to learn initiation or termination conditions, and frame the skill pretraining phase as a simple supervised learning task.
@@ 4 Skills from Unlabeled Prior data for Exploration (SUPE)
Practical implementation details. Following prior work on trajectory-segment VAEs (Ajay et al., 2021; Pertsch et al., 2021), we use a Gaussian distribution (with both mean and diagonal covariance learnable) for the trajectory encoder, the skill policy, as well as the state-dependent prior. While Pertsch et al. (2021) use a KL constraint between the high-level policy and the state-dependent prior, we use a simpler design without the KL constraint that works much better (as we show in Appendix F). To achieve this, we adapt the policy parameterization from (Haarnoja et al., 2018), where the action value is enforced to be between $-1$ and $1$ using a $\mathrm{tanh}$ transformation, and entropy regularization is applied on the squashed space. We use this policy parameterization for the high-level policy $\pi(z|s)$ to predict the skill action in the squashed space $z_{\mathrm{sqaushed}}$. We then recover the actual skill action vector by unsquashing it according to $z=\mathrm{arctanh}(z_{\mathrm{sqaushed}})$, so that it can be used by our skill policy $\pi_{\theta}(a|s,z)$. For upper-confidence bound (optimistic) estimation of the reward, ($r_{\mathrm{UCB}}(s_{0},\hat{z})$), we directly borrow the UCB estimation implementation in Li et al. (2024) (Section 3, practical implementation section in their paper), where they use a combination of the random network distillation (RND) (Burda et al., 2018) reward bonus and the predicted reward from a reward model (see Appendix C, Ours for more details). For the off-policy high-level agent, we follow Li et al. (2024) to use RLPD (Ball et al., 2023) that takes a balanced number of samples from the prior data and the online replay buffer for agent optimization. In addition to using the optimistic offline reward label, we also find that adding the RND reward bonus to the online batch is also helpful to encourage online exploration, so we use it in all our experiments.
(b) b) HumanoidMaze: medium, large, and giant.
(c) c) AntSoccer: arena and medium
(d) d) Kitchen
(e) e) Cube
(f) f) Scene
(g) g) Visual AntMaze
Figure 2: We experiment on 7 challenging, sparse-reward domains. a): AntMaze with three different layouts and the corresponding four goal locations (denoted as the red dots); b): HumanoidMaze with three layouts; c): AntSoccer with two layouts d): Kitchen; e): Cube; f): Scene; g): Visual AntMaze with colored floor and local $64\times 64$ image observations.
@@ 5 Experimental Results
    How sensitive is the performance of our method is to its hyperparameters?
@@ 5.1 Experimental setup
We conduct our experiments on 7 challenging sparse-reward domains (Figure 2, (a) - (g)). We provide a brief description of each of the domains below with more details available in Appendix D.
State-based locomotion domains: AntMaze (a), HumanoidMaze (b), AntSoccer (c). The first set of domains involve controlling and navigating a robotic agent in a complex environment. AntMaze is a standard benchmark for offline-to-online RL from D4RL (Fu et al., 2020). HumanoidMaze and AntSoccer are locomotion domains from OGBench, a offline goal-conditioned RL benchmark (Park et al., 2024a). The goal of agent (either ant or humanoid) is to navigate to a goal location in a fixed maze layout. Each of the AntMaze and HumanoidMaze domains has three different maze layouts. For AntMaze, we test on four different goal locations for each maze layout. For HumanoidMaze and AntSoccer we test on one goal, and for HumanoidMaze we use both the navigate and stitch datasets. In the AntSoccer task, the agent also needs to dribble a soccer ball to the goal location.
State-based manipulation domains: Kitchen (d), Cube (e), Scene (f). Next, we consider a set of manipulation domains that require a wide range of manipulation skills. Kitchen is a standard benchmark from D4RL where a robotic arm needs to complete a set of manipulation tasks (e.g., turn on the microwave, move the kettle) in sequence in a kitchen scene. Cube and Scene are two offline goal-conditioned RL benchmark domains from OGBench (Park et al., 2024a). For Cube, the robotic arm must arrange one or more cube objects to desired goal locations (e.g., stacking on top of each other) that mainly involves pick and place motions. For Scene, the robotic arm can interact with a more diverse set of objects: a window, a drawer, a cube and two locks that control the window and the drawer. The tasks in Scene are also longer, and require the composition of multiple atomic behaviors (e.g., locking and unlocking, opening the drawer/window, moving the cube).
In addition to the six state-based domains, we also experiment with a challenging visual-domain introduced by Park et al. (2023a), Visual AntMaze (g), where the agent has all of the proprioceptive information from the state-based AntMaze environment except for the agent’s $(x,y)$ position, which the agent must obtain from a $64\times 64$ image observation of its surroundings.
To evaluate our method on these domains, we simply take the datasets in these benchmarks and remove the reward label. We also remove any information in the transition that may reveal the information about the termination of an episode. For all of these tasks, we use a $-1$/00 sparse reward function where the agent receives $-1$ when it has not found the goal and it receives 00 when it reaches the goal location. For all of the domains above, we use the normalized return, a standard metric for D4RL (Fu et al., 2020) environments, as the main evaluation metric. For the Kitchen domain, the normalized return represents the average percentage of the tasks that are solved. For tasks in other domains, the normalized return represents the average task success rate. For all our figures in the main body, the shaded area indicates the 95% stratified bootstrap confidence interval and the solid line indicates the interquartile mean. For all plots in the appendix, the shaded area is standard error and the solid line is the mean.
@@ 5.2 Comparisons
While there is no existing method in our setting that utilizes unlabeled prior data in both the pretraining phase and the online learning phase, there are methods that use the prior data in either phase. We first consider two baselines that do not use pretraining and directly perform online learning.
Online. This baseline discards the offline data and the exploration is done with online reward bonus implemented by random network distillation (RND) (Burda et al., 2018). For all the baselines below (including our method), we add online RND bonus to the replay buffer to encourage exploration.
ExPLORe (Li et al., 2024). This baseline is similar to our method in the sense that it also uses exploration bonus and offline data to encourage exploration. The one crucial difference is that it does not perform unsupervised skill pretraining and learns a 1-step policy directly online. As we will show, pretraining is crucial for our method to find goals faster and lead to more efficient online learning. It is worth noting that the original ExPLORe method does not make use of online RND. To make the comparison fair, we additionally add online RND bonus to this baseline to help it explore better online. For completeness, we include the performance of the original ExPLORe in Figure 11.
@@
Online with Skills. We also consider two skill-based baselines where the prior data is discarded in the online phase and the high-level policy is trained from scratch online with exploration bonus. We experiment with two types of pretraining skills. The first one, is the trajectory VAE skill used in our method. The second one is from a recently proposed unsupervised offline skill discovery method where skills are pretrained to be able to traverse a learned Hilbert representation space (Park et al., 2024b) (HILP). We use the exact same high-level RL agent as our method except that the agent no longer makes use of the prior data online. It is worth noting that the baseline that uses the trajectory VAE skill is very similar to SPiRL (Pertsch et al., 2021), a prior skill-based online RL method that also pretrains skills with a trajectory VAE. The only difference is that we make two additional improvements on top of SPiRL. The first improvement is replacing the KL constraint with entropy regularization (same as our method as described in Section 4, practical implementation details). The second improvement is the online RND bonus that is also added to all other methods.
Finally, we introduce a novel baseline that also uses prior data during pretraining and online exploration, but uses HILP skills rather than trajectory-based skills.
@@
For the visual antmaze environment, we use the same image encoder used in RLPD (Ball et al., 2023). We also follow one of our baselines, ExPLORe (Li et al., 2024), to use ICVF (Ghosh et al., 2023), a method that uses task-agnostic value functions to learn image/state representations from passive data. ICVF takes in an offline unlabeled trajectory dataset with image observations and pretrain an image encoder in an unsupervised manner. Following ExPLORe, we take the weights of the image encoder from ICVF pretraining to initialize the image encoder’s weights in the RND network. To make the comparison fair, we also apply ICVF to all baselines (details in Appendix E).
Figure 3: Aggregated normalized return across seven different domains (Single-Cube and Double-Cube are two sub-domains of Cube). Ours achieves the best performance on all domains except on Scene where HILP w/ Offline Data achieves better performance. HILP w/ Offline Data is a novel baseline that we introduce which also uses the offline data both during offline skill pre-training and online learning. Section 5.2 contains details on the baselines we compare with. We omit the Online baseline on the harder domains (Cube, Scene, HumanoidMaze, and AntSoccer) as it is consistently worst than other methods. For Kitchen, we use 16 seeds. For all other domains, we use 8 seeds. All of the aggregated plots use the interquartile mean (IQM) metric following Agarwal et al. (2021) with 95% stratified bootstrap confidence intervals.